OpenStudy (anonymous):

\frac{4x^{3}+5x^{2}+3x}{x}

OpenStudy (anonymous):

\[\dfrac{4x^{3}+5x^{2}+3x}{x}\]

OpenStudy (anonymous):

Divide each term by x and add.

OpenStudy (anonymous):

HOw do you divide by x?

OpenStudy (turingtest):

\[\frac{x^a}{x^b}=x^{a-b}\]and don't forget that\[x=x^1\]

OpenStudy (anonymous):

Substitute it as 1?

OpenStudy (turingtest):

yup

OpenStudy (anonymous):

That makes it easier

OpenStudy (anonymous):

So, the answer will only have 1 exponent and it's 1?

OpenStudy (anonymous):

No, that's wrong..

OpenStudy (turingtest):

as eliassab said "divide \(each\) term by x" you can do that using the rule I gave you

OpenStudy (anonymous):

I don't know what to do in problems like these where there is just an x.

OpenStudy (turingtest):

\[x=x^1\]so for example\[\frac{x^5}{x}=\frac{x^5}{x^1}=x^{5-1}=x^4\]

OpenStudy (anonymous):

If you subtract the exponents on the top it'd be 3-2

OpenStudy (turingtest):

but the exponent on the bottom is not a 2...

OpenStudy (anonymous):

it's just an x.

OpenStudy (turingtest):

read my above posts\[x=x^1\]

OpenStudy (anonymous):

So I'd say 3-2 (-1)

OpenStudy (turingtest):

where did -2 come from?

OpenStudy (anonymous):

3-1 - 2

OpenStudy (anonymous):

3-1 = 2*

OpenStudy (turingtest):

ah... much better :) what about the other terms

OpenStudy (anonymous):

2-1 = 1.

OpenStudy (anonymous):

So, can 1 be an exponent or just leave it blank?

OpenStudy (anonymous):

and how would you divide 3x by x?

OpenStudy (turingtest):

those are the exponents and yes, you don't have to write the "to the 1" part the coefficients stay, only the x's are affected here

OpenStudy (turingtest):

for instance\[\frac{5x^4}x=5x^3\]

OpenStudy (anonymous):

That's confusing me. It's just 3x / x.

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

Or does it stay the same?

OpenStudy (turingtest):

the coefficient does not change is all I'm saying

OpenStudy (anonymous):

The answer would be 4x^2 + 5x + 3 correct?

OpenStudy (turingtest):

correct :)

OpenStudy (anonymous):

Could ya help with a couple more?

OpenStudy (anonymous):

Good Job! @careless850

OpenStudy (turingtest):

I'm sort of in an out here, and you are supposed to post each q separately, but I'll try to help when I have the chance

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks @rebeccaskell94